i filled pages to find the answer but i dont think its right please help!
Here is the problem:
How long of a shadow would light 30 feet off the ground make when shining on an object 10 feet tall that is 20 feet away from the light?

Answer by scott8148(6628)

picture a right triangle with the right angle resting on the ground

the length of the vertical side of the right angle is 30ft

20ft along the horizontal side (the side on the ground) is the 10 foot object

some distance x (the length of the shadow) beyond the 10 foot object, is the end of the horizontal side

the smaller triangle, with the 10ft vertical side and the x horizontal side is similar to
__ the larger triangle, with the 30ft vertical side and the 20+x horizontal side

10/30=x/(20+x) __ "cross" multiplying __ 200+10x=30x __ 200=20x __ 10=x